OpenStudy (anonymous):

If k(x) = 5x – 6, which expression is equivalent to (k + k)(4)? 5(4 + 4) – 6 5(5(4) – 6) – 6 54 – 6 + 54 – 6 5(4) – 6 + 5(4) – 6

OpenStudy (haseeb96):

5(4+4)-6

OpenStudy (nincompoop):

if k(x) = 3x + 2 (k + k)(4) is = ? first let us try to solve when x =4 k(4) = 3(4) + 2 = 14 for intellectual exercise, what would it mean to say 1/k (x) then? it means 1/(3x +2) expanding from this knowledge- (k+k)(x) means k(x) + k(x), which then (3x + 2) + (3x + 2) (k + k)(4) would be processed as (3(4) + 2)+(3(4) + 2) 12 + 2 + 12 + 2 = 28 go ahead and try to solve your problem

OpenStudy (nincompoop):

@Haseeb96 please do not give out answers